OpenStudy (anonymous):

Find local extrema and saddles of f(x,y)=x^3-3xy+y^3? Need help

OpenStudy (turingtest):

have you solved \(\nabla f=0\) yet?"

OpenStudy (anonymous):

yep you get <3x^2-3y,-3x+3y^2>

OpenStudy (anonymous):

then when you set it equal to zero you get y=0 and x=0 right?

OpenStudy (turingtest):

there are two solutions, that is one

OpenStudy (anonymous):

how do you find the other one?

OpenStudy (turingtest):

\[-3x+3y^2=0\implies x^2=y^4\]\[3y^4-3y=0\implies y(y^3-1)=0\implies y=\{0,1\},~x=\{0,1\}\]

OpenStudy (anonymous):

?

OpenStudy (turingtest):

solve the second equation for x^2 and substitute into the first

OpenStudy (anonymous):

so your critical points are (0,0) and (1,1)?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

so what do you do with those?

OpenStudy (turingtest):

find D for each\[D=f_{xx}(a,b)f_{yy}(a,b)-[f_{xy}(a,b)]^2\]

OpenStudy (anonymous):

find second derivatives then plug in the point values to see if they are greater or less than zero?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

so you get -9 and 27 so would (0,0) be a local min?